OpenStudy (anonymous):

√5(10-4√2) A. 14 B. 15√2 C. 5√2-4√10 D. none of the above

OpenStudy (anonymous):

D

OpenStudy (anonymous):

\[5\sqrt{10}-4\sqrt{10}\]\[\sqrt{10}\] and thats not one of the choices :)

OpenStudy (anonymous):

@sammyrayepatton helpful??:)

OpenStudy (anonymous):

yes1 thank you so much!<3

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

fan me so u can msg for more questions :D